# Reduced Order Hysteretic Magnetization Model for Composite Superconductors

### v2

@@ Abstract
In this paper, we propose the Reduced Order Hysteretic Magnetization (ROHM) model to describe the magnetization and instantaneous power loss of composite superconductors subject to time-varying magnetic fields. Once the parameters of the ROHM model are fixed based on reference simulations, it allows to directly compute the macroscopic response of composite superconductors without the need to solve the detailed current density distribution. It can then be used as part of a homogenization method in large-scale superconducting models to significantly reduce the computational effort compared to detailed simulations. In this contribution, we focus on the case of a strand with twisted superconducting filaments subject to a time-varying transverse magnetic field. We propose two levels of ROHM models: (i) a rate-independent model that reproduces hysteresis in the filaments, and (ii) a rate-dependent model that generalizes the first level by also reproducing dynamic effects due to coupling and eddy currents. We describe the implementation and inclusion of these models in a finite element framework, discuss how to deduce the model parameters, and finally demonstrate the capabilities of the approach in terms of accuracy and efficiency over a wide range of excitation frequencies and amplitudes.
  • October 25, 2024
@@
*                                    This work has been submitted to a journal for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
[…]
*                                    This work has been submitted to a journal for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
Large-scale superconducting magnets are multi-scale structures made of a high number of wound superconducting cables, which themselves consist of multifilamentary strands or complex arrangements of tapes. The magnetic response of these systems is hysteretic and rate-dependent. Hysteresis is created by the irreversible motion of flux vortices among pinning centers in superconducting elements [1, 2, 3, 4], whereas rate-dependency is caused by eddy and coupling current dynamics [5, 6, 7].
Hysteresis, eddy current, and coupling current effects induce loss in transient regimes [5] as well as field distortions [8]. Their accurate description and computation is therefore crucial, e.g., for computing the load on the cryogenic system, the temperature and stability margins, field errors, or for the design of quench protection devices.
@@
Besides these advantages, the power law introduces a strong nonlinearity in the equations, which makes the numerical models computationally demanding to solve, already for single strands or cables [11]. As a consequence, modelling large-scale superconducting systems in all their details down to the superconductor level is completely unrealistic in practice. A faster and more efficient method is necessary.
Homogenization methods which reproduce the macroscopic effects of coupling dynamics without modelling them explicitly are good candidates. The general idea of these methods is to describe the systems in terms of the average fields [12], in order to reproduce the magnetization and loss accurately, without having to solve for the detailed current density distribution, which has the potential to strongly reduce the computational cost of the simulations.
Such techniques have been extensively studied in non-superconducting electromagnetic systems, such as periodical structures with bundles of wires of any shape [13, 14, 15], described in the frequency or time domain [16, 17]. They rely on the design of equivalent conductivity and permeability parameters, that must be fitted on reference solutions.
Following a similar approach for superconducting systems requires the use of dedicated history-dependent parameters in order to reproduce their hysteretic response.
@@
The energy-based model offers several advantages. First, it gives a direct access to instantaneous dissipated and stored power by conveniently separating the magnetic field into several contributions. Second, it offers a high flexibility and modularity thanks to an approach involving several sub-elements that are combined to form the final model. Third, it is consistently defined as a vector, and not scalar model from the beginning [27], and hence applies to two-dimensional (2D) or three-dimensional (3D) settings. Fourth, it can be directly generalized to include rate-dependent effects [28]. Finally, the equations of this model consist in simple explicit tests, which leads to straightforward implementations and very efficient simulations.
In this paper, we adapt and extend the state-of-the-art energy-based model to composite superconductor hysteresis by focussing on the case of a superconducting multifilamentary strand subject to a transverse magnetic field, i.e., a field perpendicular to the strand axis. This situation is relevant for most applications and constitutes a first step towards the homogenization of a complete magnet winding. This is also a challenging problem since, in addition to the filament hysteresis effect, it also involves two types of rate-dependent effects, eddy currents and interfilamentary coupling currents in the strand matrix, as well as filament hysteresis change due to coupling and eddy currents. For this paper, we assume that the strand carries no transport current. The inclusion of transport current will be the focus of further work. We also consider a constant temperature and do not model thermal effects.
We propose the Reduced Order Hysteretic Magnetization (ROHM) model, and present two variations of it: (i) a rate-independent model which describes the superconducting hysteresis, and (ii) a rate-dependent model, which generalizes the former by also describing eddy current and coupling current effects, in addition to the superconducting hysteresis.
The ROHM model is general and can be adapted to a variety of superconductors or composite superconductors, such as strands, tapes, stacks of tapes, cables, or bulks. It can also be easily implemented in any FE framework that allows for user-defined material properties.
This paper is structured as follows. In Section 2, we describe the response of a multifilamentary strand in terms of magnetization and loss for transverse fields with a range of frequencies and amplitudes, in order to motivate the design of the ROHM model. In Section 3, we present the equations defining the rate-independent and rate-dependent ROHM models. In Section 4, we discuss the implementation of these equations and we describe how to include them in a 2D FE framework with a $\phi$-formulation. We then propose in Section 5 a parameter identification approach for fixing the ROHM model parameters based on reference solutions. Finally, in Sections 6 and 7, we apply the ROHM model in rate-independent and rate-dependent situations, respectively, and demonstrate its potential in the context of homogenization techniques.
@@ 2 Multifilamentary Strand Dynamics
In this section, we describe the response of a composite superconducting strand subject to a transverse magnetic field in terms of power loss and magnetization. To this end, we model the detailed current density distribution inside the strand with a FE method, the CATI method [30] implemented in FiQuS [31, 32, 33], and we analyze the different dynamics resulting from the strand composite structure.
@@ 2.3.1 Rate-independent regime
At low frequencies (here for $f\lesssim 0.01$ Hz), the rate of change of the transverse magnetic field generates negligible coupling and eddy currents in the matrix. In this regime, the superconducting filaments behave almost uncoupled and most of the loss is due to hysteresis in them. A typical magnetization loop is shown in Fig. 2. The magnetization decreases when the field increases as a result of the field-dependent critical current density.
@@ 3 Reduced Order Hysteretic Magnetization Model
The simulation time for each of the simulations presented in the previous section is of the order of one hour with the CATI method [30]. This is fast and convenient enough to get a good understanding of a single strand response with a detailed description of the current density distribution (as discussed in [30] or shown later in Fig. 27). However, to simulate transients in a full-scale magnet containing thousands of strands, one cannot afford models involving such a detailed description at the strand level.
Instead, one can use an alternative model that produces accurate macroscopic response in terms of power loss and magnetization, which are both history-dependent and rate-dependent, but that does not rely on a detailed calculation of the current density. We present such a model in this section: the Reduced Order Hysteretic Magnetization (ROHM) model. This model is adapted from the energy-based model for ferromagnetic materials [27, 28].
We start in Section 3.1 by clarifying how the ROHM model can be used to model the strand. We then define the building blocks, referred to as cells, that will be combined for the construction of the ROHM model. Each cell defines a local hysteretic relationship between the magnetic field $\boldsymbol{h}$ and the magnetic flux density $\boldsymbol{b}$. We present in Section 3.2 a rate-independent cell, that we call the superconductor cell (S cell). In Section 3.3, we add contributions to account for eddy current and coupling current effects, resulting in a rate-dependent cell, that we call the composite superconductor cell (CS cell). Finally, we show in Section 3.4 how to combine several cells into a chain of cells, defining a complete ROHM model.
@@ 3.3 Composite superconductor cell - CS cell
For regimes in which the superconducting strand response is strongly rate-independent, the S cell presented in the previous section must be generalized. Starting from the S cell, we propose to account for the magnetization due to eddy currents and coupling currents by adding two new contributions in the magnetic field decomposition, $\boldsymbol{h}_{\text{eddy}}$ and $\boldsymbol{h}_{\text{coupling}}$, such that Eq. (8) is rewritten
@@ 4.1 S cell
$\boldsymbol{h}_{\text{rev}}$ =$&\boldsymbol{h}_{\text{rev(p)}},&&\text{if }\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|\leq\kappa,\\ &\boldsymbol{h}-\kappa\frac{\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}}{\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|},&&\text{otherwise,}\end{aligned}\right.$
$=\boldsymbol{\mathcal{U}}_{\kappa}(\boldsymbol{h},\boldsymbol{h}_{\text{rev(p)}}).$ (27)
@@
If the irreversibility parameter $\kappa$ is not constant but a function of $\boldsymbol{b}$ to account for field-dependent $j_{\text{c}}$, the update rule is no longer explicit. In such a case, an approximate solution can be computed easily with a few fixed point iterations, that is, by solving successively Eq. (4.1) with updated values of $\kappa(\boldsymbol{b})=\kappa(\mu_{0}\boldsymbol{h}_{\text{rev}})$ until $\boldsymbol{h}_{\text{rev}}$ does no longer change significantly.
The implementation of the S cell is very simple, as it entirely consists in the test of Eq. (4.1). It only requires the knowledge of the previous reversible field $\boldsymbol{h}_{\text{rev}}$, which has to be stored as an internal variable during the numerical simulation.
Note that by contrast with ferromagnetic hysteresis with a nonlinear saturation law [27, 29], in this context of superconducting hysteresis, the relationship Eq. (9) is linear. The update rule Eq. (4.1) is therefore exact [29], up to the time discretization error.
@@ 4.2 CS cell
For the CS cell described in Section 3.3, the approach is similar. However, two tests are now necessary to account for the two dry friction elements. With
$\boldsymbol{g}=\boldsymbol{h}_{\text{rev}}+\boldsymbol{h}_{\text{eddy}}+\boldsymbol{h}_{\text{coupling}},$ (28)
and $\boldsymbol{g}_{\text{(p)}}$ its value at the previous time step. We start by updating $\boldsymbol{g}$, as a function of the new $\boldsymbol{h}$ as follows:
$\boldsymbol{g}=\boldsymbol{\mathcal{U}}_{\kappa}(\boldsymbol{h},\boldsymbol{g}_{\text{(p)}}),$ (29)
with the update rule defined by Eq. (4.1) but as a function of $\boldsymbol{g}_{\text{(p)}}$ instead of $\boldsymbol{h}_{\text{rev(p)}}$. This operation contains the first test, after which we can already compute the irreversible field $\boldsymbol{h}_{\text{irr}}=\boldsymbol{h}-\boldsymbol{g}$.
@@ 4.4 Inclusion in a finite element $\phi$-formulation
As the hysteresis law is driven by the magnetic field $\boldsymbol{h}$, it is easier to consider an $h$-conform formulation, such as the $h$-$\phi$-formulation [39, 40]. In this particular case with no net currents in $\Omega_{\text{m}}$ and a non-conducting $\Omega_{\text{m}}^{\text{C}}$, this formulation can be reduced to a $\phi$-formulation [29].
Starting from an initial solution, it consists in finding a magnetic field $\boldsymbol{h}=\textrm{{grad}}\,\phi$ with $\phi\in\Phi(\Omega)$ such that, at subsequent time instants, $\forall\boldsymbol{h}^{\prime}=\textrm{{grad}}\,\phi^{\prime}$ with $\phi^{\prime}\in\Phi_{0}(\Omega)$,
@@
where the notation $\left(\boldsymbol{v}\,,\boldsymbol{w}\right)_{\Omega}$ denotes the integral over $\Omega$ of the dot product of any two vector fields $\boldsymbol{v}$ and $\boldsymbol{w}$. The space $\Phi(\Omega)$ contains functions in $H^{1}(\Omega)$ that satisfy appropriate essential boundary conditions. The space $\Phi_{0}(\Omega)$ is the same space but with homogeneous essential boundary conditions. These function spaces are discretized with lowest-order node functions [41].
@@ 6.1 Parameter identification
Figure 18: Comparison of the ROHM model results with the reference solution. Unidirectional transverse magnetic field. Chain of S cells with $N=6$ cells, parameters of Fig. 17. (a) Harmonic excitation $h_{\text{app}}^{(1)}(t)$. (b) Biharmonic excitation $h_{\text{app}}^{(2)}(t)$. (Up) Internal magnetic field, the input of the ROHM model. (Middle) Magnetization loop. The curves of the ROHM model are colored as a function of the instantaneous power, see the bottom plots for the legend. (Down) Dissipated power per unit length as a function of time, equal to $a\,p_{\text{irr}}(t)$, with $a=\pi d^{2}/4$ the surface area of the strand cross section.

### v3

@@ Abstract
In this paper, we propose the Reduced Order Hysteretic Magnetization (ROHM) model to describe the magnetization and instantaneous power loss of composite superconductors subject to time-varying magnetic fields. Once the parameters of the ROHM model are fixed based on reference simulations, it allows to directly compute the macroscopic response of composite superconductors without the need to solve the detailed current density distribution. It can then be used as part of a homogenization method in large-scale superconducting models to significantly reduce the computational effort compared to detailed simulations. In this contribution, we focus on the case of a strand with twisted superconducting filaments subject to a time-varying transverse magnetic field. We propose two variations of the ROHM model: (i) a rate-independent model that reproduces hysteresis in the filaments, and (ii) a rate-dependent model that generalizes the first level by also reproducing dynamic effects due to coupling and eddy currents. We then describe the implementation and inclusion of the ROHM model in a finite element framework, discuss how to deduce the model parameters, and finally demonstrate the capabilities of the approach in terms of accuracy and efficiency over a wide range of excitation frequencies and amplitudes.
  • February 5, 2025
@@ 1 Introduction
Large-scale superconducting magnets are multi-scale structures made of multi-turn windings of superconducting cables, which themselves consist of multifilamentary strands or complex arrangements of tapes. The magnetic response of these systems is hysteretic and rate-dependent because of small-scale current density dynamics. Hysteresis is created by the irreversible motion of flux vortices among pinning centers in superconducting elements [1, 2, 3, 4]. Rate-dependency is caused by eddy current and coupling current dynamics [5, 6, 7].
Hysteresis, eddy current, and coupling current effects induce loss in transient regimes [5] as well as field distortions [8]. Their accurate description and computation is therefore crucial, e.g., for computing the load on the cryogenic system, temperature and stability margins, field errors, or for the design of quench protection devices.
@@
Besides these advantages, the power law introduces a strong nonlinearity in the equations, which makes numerical models of superconducting systems computationally demanding to solve, already for single strands or cables [11, 12, 13, 14]. As a consequence, modelling large-scale superconducting systems in all their details down to the superconductor level is completely unrealistic in practice. It is often necessary to group conductors together in order to lower simulation time [15, 16], and multi-scale procedures [17] or domain decomposition methods [13, 18] combined with parallel computing are being developed to speed up simulations.
Homogenization methods describing systems in terms of the average fields [19] without modelling the small-scale field distributions explicitly are also good candidates for obtaining even more efficient large-scale superconducting magnet models. The general idea of these methods consists in replacing complex conducting structures by fictitious, uniform, homogenized materials with special properties. These properties are designed in order to reproduce directly the macroscopic effects of the small-scale current density distribution, e.g., magnetization and loss, without modelling that current density distribution explicitly. Once the homogenized material properties are defined by fitting their predictions with those of reference models, they can be used to reproduce accurately the response of large-scale systems with a strongly reduced computational cost compared to conventional detailed simulations.
Such homogenization methods have been extensively studied in non-superconducting electromagnetic systems, such as periodical structures with bundles of wires of any shape [20, 21, 22], described in the frequency or time domain [23, 24]. They rely on the definition of equivalent permeability and conductivity values (or functions), to account for proximity effect and skin effect, respectively.
Following a similar approach for superconducting systems requires to define history-dependent material properties in order to reproduce a hysteretic response. This is the objective of this paper.
@@
The energy-based model offers several advantages compared to other models. First, it gives a direct access to instantaneous dissipated and stored power by conveniently separating the magnetic field into several contributions. Second, it offers a high flexibility and modularity thanks to an approach involving several sub-elements that are combined to form the final model. Third, it is consistently defined as a vector model from the beginning [34], and hence applies to two-dimensional (2D) or three-dimensional (3D) settings. Fourth, it can be directly generalized to include rate-dependent effects [35]. Finally, the equations of this model consist in simple explicit tests, which leads to straightforward implementations and very efficient simulations.
In this paper, we adapt and extend the state-of-the-art energy-based model to composite superconductor hysteresis, paying attention to preserving all the above-mentioned advantages. In this first contribution, we focus on the case of a superconducting multifilamentary strand subject to a transverse magnetic field, i.e., a field perpendicular to the strand axis, which is relevant for most magnet applications. The multifilamentary strand is also a challenging problem. Indeed, in addition to the filament hysteresis effect, it also involves two types of rate-dependent effects, interfilament coupling currents and eddy currents in the strand matrix, as well as filament hysteresis change due to coupling and eddy currents.
In this paper, we assume that the strand carries no transport current. Accounting for transport current and the computing the associated voltage is necessary for the homogenization of a complete magnet winding, but will be the focus of a further work. We also consider a constant temperature and do not model thermal effects.
[…]
Although we apply it on multifilamentary strands only, the ROHM model can be adapted to a variety of superconductors or composite superconductors, such as strands, tapes, stacks of tapes, cables, or bulks.
This paper is structured as follows. In Section 2, we describe the response of a multifilamentary strand in terms of magnetization and loss for transverse fields with a range of frequencies and amplitudes, in order to motivate the design of the ROHM model. In Section 3, we present the equations defining the rate-independent and rate-dependent variations of the ROHM model. In Section 4, we discuss the implementation of these equations and we describe how to include them in a 2D FE framework with a $\phi$-formulation. We then propose in Section 5 a parameter identification approach for fixing the ROHM model parameters based on reference solutions. Finally, in Sections 6 and 7, we apply the ROHM model in rate-independent and rate-dependent situations, respectively, and demonstrate its potential as a homogenized model.
@@ 2 Multifilamentary Strand Dynamics
In this section, we describe the response of a composite superconducting strand subject to a transverse magnetic field in terms of the induced power loss and magnetization. To this end, we model the detailed current density distribution inside the strand with a FE method, the CATI method [37] implemented in FiQuS [38, 39, 40], and we analyze the different dynamics resulting from the strand composite structure.
@@ 2.3.1 Rate-independent regime
At low frequencies (here for $f\lesssim 0.01$ Hz), the rate of change of the transverse magnetic field generates negligible coupling and eddy currents in the matrix [5, 17]. In this regime, the superconducting filaments behave almost as uncoupled and most of the loss is due to hysteresis in them. A typical magnetization loop is shown in Fig. 2. The magnetization decreases when the field increases as a result of the field-dependent critical current density.
@@ 3 Reduced Order Hysteretic Magnetization Model
The simulation time for each of the simulations presented in the previous section is of the order of one hour with the CATI method [37]. This is fast and convenient enough to get a good understanding of a single strand response with a detailed description of the current density distribution (as discussed in [37] or shown later in Fig. 27). However, to simulate transients in a full-scale magnet cross section containing thousands of strands, one cannot afford models involving such a detailed description at the strand level.
Instead, one can use an alternative model that produces accurate macroscopic response in terms of power loss and magnetization, which are both history-dependent and rate-dependent, but that does not rely on a detailed calculation of the current density. We present such a model in this section: the Reduced Order Hysteretic Magnetization (ROHM) model. This model is inspired from the energy-based model for ferromagnetic materials [34, 35].
We start in Section 3.1 by clarifying how the ROHM model can be used to model a superconducting strand.
We then define the building blocks, referred to as cells, that will be used for the construction of the ROHM model. Each cell defines a local and elementary hysteretic relationship between the magnetic field $\boldsymbol{h}$ as an input, and the magnetic flux density $\boldsymbol{b}$ as an output. We present in Section 3.2 a rate-independent cell, that we call the superconductor cell (S cell). In Section 3.3, we add components to account for eddy current and coupling current effects, resulting in a rate-dependent cell, that we call the composite superconductor cell (CS cell). Inside each cell, the magnetic field vector is expressed as the sum of distinct components: 2 components for the S cell, and 4 components for the CS cell. These components describe the interplay between different magnetization mechanisms and allow for a clear separation between stored and dissipated energies.
A single cell is not sufficient for an accurate description of the strand response. We show in Section 3.4 how several cells can be combined into a chain of cells, defining the complete ROHM model.
@@ 3.3 Composite superconductor cell - CS cell
For regimes in which the superconducting strand response is strongly rate-dependent, the S cell presented in the previous section must be generalized. Starting from the S cell, we propose to account for the magnetization due to eddy currents and coupling currents by adding two new contributions in the magnetic field decomposition, $\boldsymbol{h}_{\text{eddy}}$ and $\boldsymbol{h}_{\text{coupling}}$, such that Eq. (8) is rewritten
@@ 4.1 S cell
$\boldsymbol{h}_{\text{rev}}$ =$&\boldsymbol{h}_{\text{rev(p)}},&&\text{if }\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|\leq\kappa,\\ &\boldsymbol{h}-\kappa\frac{\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}}{\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|},&&\text{otherwise,}\end{aligned}\right.$ (27)
@@
If the irreversibility parameter $\kappa$ is not constant but a function of $\boldsymbol{b}$ to account for field-dependent $j_{\text{c}}$, the update rule is no longer explicit. In such a case, an approximate solution can be computed easily with a few fixed point iterations, that is, by solving successively Eq. (27) with updated values of $\kappa(\boldsymbol{b})=\kappa(\mu_{0}\boldsymbol{h}_{\text{rev}})$ until $\boldsymbol{h}_{\text{rev}}$ does no longer change significantly.
The implementation of the S cell is very simple, as it entirely consists in the test of Eq. (27). It only requires the knowledge of the previous reversible field $\boldsymbol{h}_{\text{rev}}$, which has to be stored as an internal variable during the numerical simulation.
Note that by contrast with ferromagnetic hysteresis with a nonlinear saturation law [34, 36], in this context of superconducting hysteresis, the relationship Eq. (9) is linear. The update rule Eq. (27) is therefore exact [36], up to the time discretization error.
@@ 4.2 CS cell
For the CS cell described in Section 3.3, the approach is similar. However, two tests are now necessary to account for the two dry friction elements.
We start by updating $\boldsymbol{g}=\boldsymbol{h}_{\text{rev}}+\boldsymbol{h}_{\text{eddy}}+\boldsymbol{h}_{\text{coupling}}$ (see Fig. 10). Denoting $\boldsymbol{g}_{\text{(p)}}$ its value at the previous time step, we define the update rule for $\boldsymbol{g}$ as follows:
$\boldsymbol{g}$ =$&\boldsymbol{g}_{\text{(p)}},\text{ if }\|\boldsymbol{h}-\boldsymbol{g}_{\text{(p)}}\|\leq\kappa\text{ and }\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|\leq\kappa,\\ &\boldsymbol{h}-\kappa\frac{\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}}{\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|},\text{ otherwise,}\end{aligned}\right.$ (28)
where the second necessary condition in the first case ensures that $\|\boldsymbol{h}_{\text{irr}}\|<\kappa$ is possible only if $\dot{\boldsymbol{b}}=\boldsymbol{0}$. From $\boldsymbol{g}$, we can compute $\boldsymbol{h}_{\text{irr}}=\boldsymbol{h}-\boldsymbol{g}$.
@@ 4.4 Inclusion in a finite element $\phi$-formulation
As the hysteresis law is driven by the magnetic field $\boldsymbol{h}$, it is easier to consider an $h$-conform formulation, such as the $h$-$\phi$-formulation [46, 47]. In this particular case with no net currents in $\Omega_{\text{m}}$ and a non-conducting $\Omega_{\text{m}}^{\text{C}}$, the $h$-$\phi$-formulation can be reduced to a $\phi$-formulation [36]. The inclusion of the ROHM model in a $b$-conform formulation is less straightforward and is discussed in [48].
Starting from an initial solution, the $\phi$-formulation consists in finding a magnetic field $\boldsymbol{h}=\textrm{{grad}}\,\phi$ with $\phi\in\Phi(\Omega)$ such that, at subsequent time instants, $\forall\boldsymbol{h}^{\prime}=\textrm{{grad}}\,\phi^{\prime}$ with $\phi^{\prime}\in\Phi_{0}(\Omega)$,
@@
where the notation $\left(\boldsymbol{v}\,,\boldsymbol{w}\right)_{\Omega}$ denotes the integral over $\Omega$ of the dot product of any two vector fields $\boldsymbol{v}$ and $\boldsymbol{w}$. The function space $\Phi(\Omega)$ contains scalar fields in $H^{1}(\Omega)$ that satisfy appropriate essential boundary conditions. The function space $\Phi_{0}(\Omega)$ is the same function space but with homogeneous essential boundary conditions. These function spaces are discretized with lowest-order node functions [49].
@@ 6.1 Parameter identification
Figure 18: Comparison of the ROHM model results with reference solutions. Unidirectional transverse magnetic field. Chain of S cells with $N=6$ cells, parameters of Fig. 17. (a) Harmonic excitation $h_{\text{app}}^{(1)}(t)$. (b) Biharmonic excitation $h_{\text{app}}^{(2)}(t)$. (Up) Internal magnetic field, the input of the ROHM model. (Middle) Magnetization loop. The curves of the ROHM model are colored as a function of the instantaneous power, according to the values in the bottom plots. (Down) Dissipated power per unit length as a function of time, equal to $a\,p_{\text{irr}}(t)$, with $a=\pi d^{2}/4$ the surface area of the strand cross section.